OpenStudy (anonymous):

Express the function in the form (f ºgºh)(x) and identify f(x), g(x) and h(x) The function is : H(x)=1- (3^(x^2))

OpenStudy (anonymous):

fºgºh means : f(g(h(x)))

OpenStudy (anonymous):

you could use \[h(x)=x^2\] \[g(x)=3^x\] \[f(x)=1-x\]

OpenStudy (anonymous):

take a number for x and what do you do? 1) square it 2) raise three to that power 3) subtract the result from 1 that explains the answer i wrote i hope

OpenStudy (anonymous):

:P thanks, ill try it.

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

WOW~ IT WORKS DUDE!!! :p thanks!! could you check my other one,